# An efficient wavelet-based physics-informed neural networks for singularly perturbed problems

### v1

@@ 3 Results and discussions
Predator-prey interaction model [51]:
A predator-prey system is an interconnected population of two species in which one species, the predator, actively hunts and consumes the other species, the prey. A singular perturbation argument detects limit cycle behavior and describes other dynamical situations observed in the predator-prey interaction. The Predator-prey interaction model is mathematically defined as follows:
$\begin{cases}\dfrac{dx}{dt}=kxy-\phi_{1}(x),\\[6.0pt] \dfrac{dy}{dt}=\phi_{2}(y)-\dfrac{1}{\epsilon}xy,\\ \end{cases}$ (10)
where $x(t)\geq 0$ denotes the predator density, $y(t)\geq 0$ represents the prey density, $\phi_{1}(x)$ denotes the natural death rate of the predator population, and $\phi_{2}(y)$ denotes the change in prey population due to natural growth. Here, we consider $\phi_{1}(x)=dx$, which denotes the exponential decay rate of the predator population due to natural death. Moreover, $\phi_{2}(y)=ry\Bigl{(}1-\dfrac{y}{m}\Bigr{)}$, with $r>0$, $m>0$ is a logistic growth rate for the prey population. The parameter $k$ is referred to as the conversion rate.
For the numerical computations, we set $k=0.5,~{}d=1,~{}r=1,~{}m=0.1,$ and $\epsilon=2^{-8}$ (singular perturbation parameter), let the initial density of both predator and prey be same, $x(0)=1$ and $y(0)=1$. The dynamics of the system are computed till $t=1$.
We consider parameters for this problem 3.6 that are presented in Table 12. The results for this example are discussed in Table 12 and Figure 12. Specifically, Table 12 displays a comparison between the W-PINN, the traditional PINN, and the PINN described in [41] that employs a Gaussian wavelet as an activation function for $\epsilon=2^{-8}$. It is evident from this table that the W-PINN with a Gaussian wavelet addresses the singularity more effectively. Further, the large errors in the $x$ and $y$ solutions obtained with the PINN using a sigmoid activation function highlight its failure to detect the singularity. While the PINN with the Gaussian wavelet as an activation function accurately approximates the solution $x$ without any singularity, it fails to satisfactorily resolve the singularity in the solution $y$. On the other hand, the W-PINN exhibits lower error and effectively handles the singularity in the $y$ solutions. These observations can also be drawn from Figure 12, in which solutions using PINN and W-PINN are plotted for $\epsilon=2^{-8}$.
@@ 4 Conclusion
This paper successfully demonstrated the advantages of W-PINNs over traditional PINNs and PINNs with wavelet activation functions for addressing singularly perturbed problems. The wavelet theory served as the foundation for the development of W-PINN, which combined the learning efficiency of PINN with the localization properties of wavelets (both in scale and space) to capture nonlinear information within localized regions. A limitation of the proposed method is that it will perform adequately when the resolution of the wavelet family is chosen over a specific small interval. A natural extension of this work involves applying it to singularly perturbed partial differential equations. W-PINNs may also prove effective for other problems that exhibit singular behavior, such as fractional differential models, which inherently possess such characteristics.
@@ Acknowledgments
Ratikanta Behera acknowledges the support of the Science and Engineering Research Board (SERB), Department of Science and Technology, India, under Grant No. EEQ/2022/001065.

### v2

@@ 3 Results and discussions
A heat conduction problem with large gradients [44]:
The heat conduction problem investigates temperature flow when an intense heat source suddenly appears. It represents a practical challenge often seen in fusion applications, where researchers need to understand and model rapid heat transfer.
Table 10: Parameters used for Example 3.5.
Parameters Value
Set of resolutions $(J_{x},J_{t})$ [-3,5], [-3,5]
Number of hidden layers 6
Neurons per layer 50
Number of collocation points $10^{4}$
Number of boundary points $10^{3}$
Number of initial points $500$
Table 11: Comparison of various methods for Example 3.5 with $\epsilon=0.15$.
Methods $L_{2}$-error Average Training Time
Conventional PINN $1.07$0.11 ×$\times 10^{0}$
SA-PINN [24] 1.76 $\times$ $10^{-3}$ 66.75 min
MMPINN-DNN [23] $5.01$1.52 ×$\times 10^{-4}$ 11.02 min
W-PINN (proposed method) $3.96$1.3 ×$\times 10^{-4}$ 9.5 min
The following mathematical model includes a small positive constant $\epsilon$ that creates steep temperature gradients, making it a useful test case for our developed method:
$\begin{cases}\dfrac{du}{dt}=\dfrac{d^{2}u}{dx^{2}}+f(x,t),~{}x\in(-1,1),~{}t\in[0,1],\\[6.0pt] u(x,0)=(1-x^{2})\exp\left(\dfrac{1}{1+\epsilon}\right),~{}x\in(-1,1),\\[6.0pt] u(-1,t)=0,~{}u(1,t)=0,~{}t\in(0,1],\end{cases}$ (10)
where $f$ is chosen such that $u(x,t)=(1-x^{2})\exp\left(\dfrac{1}{(2t-1)^{2}+\epsilon}\right)$ is the exact solution.
The behavior of this model exhibits interesting characteristics depending on the value of the parameter $\epsilon$. When $\epsilon$ is relatively large, the solution $u$ demonstrates smooth variations across the computational domain. However, when $\epsilon$ becomes very small, the solution shows dramatic changes near $t=0.5$, exhibiting distinct multi-scale characteristics. An important observation is the relationship between the supervised loss term and the residual term - while the boundary conditions ensure the supervised loss term remains minimal (as the boundary value is 0 and initial value stays below $\epsilon$), the residual term grows considerably as $\epsilon$ decreases. This is quantitatively demonstrated as, when $\epsilon=0.15$ the ratio of $\mathcal{L}_{\text{bc}}:\mathcal{L}_{\text{ic}}:\mathcal{L}_{\text{res}}=1:10:10^{7}.$ Traditional PINN methods generally perform well for smooth problems, but face challenges when dealing with such large disparities between supervised and residual terms. Our proposed W-PINN can effectively handle such loss imbalances.
Figure 9: The exact solution (left) and the prediction of W-PINN (right) with $\epsilon=0.15$ for Example 3.5 .
Figure 10: Point-wise absolute error for Example 3.5. From left to right: SA-PINN, MMPINN-DNN, and W-PINN.
Table 10 represents the parameters of the network for this Example 3.5. Further, Table 11 presents a comparative analysis of different methods for solving Example 3.5 with $\epsilon=0.15$, evaluating them based on $L_{2}$-error and average training time. The conventional PINN fails to get an accurate prediction. The SA-PINN [24] shows significant improvement in accuracy but requires substantial computational resources, as evident from its lengthy training time. The MMPINN-DNN [23] achieves better accuracy with reduced computational cost. However, our proposed W-PINN approach demonstrates the most promising balance between accuracy and efficiency. It achieves comparable or better accuracy than other advanced methods while reducing the computational overhead, making it particularly attractive for practical applications. This improvement in both accuracy and computational efficiency highlights the effectiveness of our proposed W-PINN in handling multi-scale characteristics.
Figure 11: Loss curve of PINN (left) and W-PINN (right) for Example 3.5 with $\epsilon=0.15$.
Figure 9 demonstrates the similarity between the exact solution and the W-PINN’s prediction for Example 3.5. Figure 10 shows point-wise absolute error comparisons across three different approaches: SA-PINN, MMPINN-DNN, and W-PINN. This error plot shows that errors for W-PINN throughout the domain remain consistently low except for extremely small regions near corners. Further, Figure 11 demonstrates the training progress of conventional PINN and W-PINN by showing their loss components over iterations. PINN shows unstable behavior with the boundary and initial condition losses initially increasing, suggesting potential training difficulties. In contrast, W-PINN exhibits a much more stable and efficient training pattern, where all three loss components (residual, initial condition, and boundary condition) consistently decrease from the start, reaching lower magnitudes in fewer iterations. This indicates that W-PINN achieves better optimization performance than the standard PINN approach.
Example 3.6.
Helmholtz equation with high-frequency:
The Helmholtz equation is a crucial elliptic partial differential equation that models electromagnetic wave behavior. It appears in wide applications in physics and engineering. The equation combines the Laplacian operator ($\Delta u$) with a wave number term ($c^{2}$), and is typically studied on bounded domains with appropriate boundary conditions. In the high-frequency regime, this equation becomes particularly challenging to solve numerically due to the oscillatory nature of its solutions.
$\begin{cases}\Delta u(x,y)+c^{2}u(x,y)=f(x,y),&(x,y)\in\Omega=(-1,1)\times(-1,1),\\ u(x,y)=p(x,y),&(x,y)\in\partial\Omega.\end{cases}$ (11)
$f$ and $p$ are obtained in such a way $u(x,y)=\sin(b_{1}\pi x)\sin(b_{2}\pi y)$ is the exact solution.
Due to the significant initial imbalance between residual and supervised components, conventional PINN fails to provide accurate predictions as the optimization process becomes heavily skewed towards residual minimization [45].
Table 12: Parameters used for Example 3.6.
Parameters Value
Set of resolutions $(J_{x},J_{t})$ [-4,5], [-4,5]
Number of hidden layers 6
Neurons per layer 50
Number of collocation points $10^{4}$
Number of boundary points $10^{3}$
Table 13: Comparison of various methods for Example 3.6.
Methods $L_{2}$-error Average Training Time
Conventional PINN $4.93$1.56 ×$\times 10^{-2}$ 23.18 min
SA-PINN [24] 1.27±1.11$\times$ $10^{-2}$ 152.88 min
MMPINN-MFF[23] $6.56$4.17 ×$\times 10^{-4}$ 22.75 min
MMPINN-INN[23] $2.13$0.43 ×$\times 10^{-4}$ 64.29 min
W-PINN (proposed method) $3.12$0.71×$\times 10^{-4}$ 6.3 min
To test the performance of W-PINN, we set the model parameters as $c=1,b_{1}=1$, and $b_{2}=8$. Table 12 represents the parameters of the network for this Example 3.6. The comparative analysis presented in Table 13 demonstrates that W-PINN achieves comparable or superior accuracy in terms of $L_{2}$-error relative to state-of-the-art methods in recent literature. Along with its significantly reduced training time, W-PINN is established as a particularly attractive framework. Figure 12 illustrates the remarkable agreement between the analytical solution and W-PINN predictions across the entire domain. In Figure 13, cross-sectional comparisons at various x-coordinates further validate the method’s accuracy, with predicted solutions being indistinguishable from the exact solutions.
Figure 12: From left to right: The exact solution, the predicted solution, and the Point-wise Error using W-PINN for the Helmholtz equation 3.6
Figure 13: Comparison of the exact solution and the prediction using W-PINN for Example 3.6.
Example 3.7.
Allen-Cahn reaction-diffusion equation:
The Allen-Cahn reaction-diffusion PDE is a widely used model in materials science, particularly for simulating phase separation processes in metallic alloys [46, 47]. The equation describes the evolution of an order parameter, $u$, which represents the phase state of the material. The PDE combines a diffusion term, with a nonlinear reaction term that drives the phase separation. The Allen-Cahn equation used in this study is defined as:
$\begin{cases}u_{t}-\epsilon u_{xx}+5u^{3}-5u=0,\quad x\in[-1,1],\quad t\in[0,1],\\ u(t,-1)=u(t,1),\\ u_{x}(t,-1)=u_{x}(t,1),\\ u(x,0)=x^{2}\cos(\pi x),\end{cases}$ (12)
here $\epsilon$ denotes a singularly perturbed parameter. For numerical computations, we choose $\epsilon=10^{-4}$. Moreover, the exact solution to this problem is unknown, so we obtained a numerical solution using Scipy solver and treated it as an exact solution. The Allen-Cahn equation is particularly challenging due to its rapid changes across space and time, making it difficult to model accurately. Additionally, its periodic boundary conditions provide an extra layer of complexity, making it an ideal benchmark for testing the model’s capabilities.
Table 14: Parameters used for Example 3.7.
Parameters Value
Resolutions $(J_{x},J_{t})$ [-5,6], [-5,5]
Number of hidden layers 6
Neurons per layer 100
Number of collocation points $2\times 10^{4}$
Number of boundary points $2\times 10^{3}$
Number of initial points $10^{3}$
Table 15: Comparison of various methods for Example 3.7 with $\epsilon=10^{-4}$.
Methods $L_{2}$-error
Conventional PINN $0.96\pm 0.06$
Time-adaptive approach [48] $8.0\times 10^{-2}$$\pm$ $0.56\times 10^{-2}$
SA-PINN [24] $2.1\times 10^{-2}\pm 1.21\times 10^{-2}$
W-PINN (proposed method) $4.8\times 10^{-2}\pm 0.6\times 10^{-2}$
Figure 14: Top: Predicted solution of Allen-Cahn equation 3.7 via W-PINN. Bottom: point-wise absolute error distribution.
Figure 15: Comparison between exact solution (solid blue line) and predicted solution (dashed orange line) at different time instances for Example 3.7.
Table 15 presents a comparative analysis of different PINN-based methods. The conventional PINN completely fails to capture the complex dynamics of the problem. While the time-adaptive method [48] shows some improvement, its accuracy remains unsatisfactory. In contrast, our proposed W-PINN achieves accuracy comparable to SA-PINN [24] while demonstrating computational efficiency with approximately twice the speed ( 30 ms/iteration) of SA-PINN. Figure 14 illustrates the predicted dynamics using the W-PINN method, where the predominance of dark blue regions in the error distribution plot indicates consistently low prediction errors across the solution domain. Further, Figure 15 presents a comparison of W-PINN prediction with the exact solution at three different time stamps, validating the model’s ability to accurately capture the temporal evolution of the system dynamics.
Example 3.8.
Maxwell’s Equation:
Maxwell’s equations represent the fundamental principles governing electromagnetic theory, describing the relationships between electric and magnetic fields and their interaction with matter. Their applications span numerous fields of engineering and physics. In telecommunications, these equations govern the propagation of electromagnetic waves through optical fibers and wireless channels. The medical industry relies heavily on Maxwell’s equations for diagnostic imaging technologies, particularly in magnetic resonance imaging (MRI) [49]. In the aerospace sector, these equations are essential for radar system design and satellite communication systems. In their differential form, Maxwell’s equations are expressed as:
[…]
where $\boldsymbol{E}$ is the electric field vector, $\boldsymbol{H}$ is the magnetic field vector. The material properties are characterized by $\mu$, the magnetic permeability, which quantifies the medium’s response to magnetic fields, and $\epsilon$, the electric permittivity, which describes the medium’s capacity to store electrical energy. The operators $\nabla$ and $\times$ represent the del operator and cross product, respectively.
A significant challenge in solving these equations arises from their inherent rapid oscillations in both space and time. This characteristic poses particular difficulties for traditional PINNs, especially when dealing with heterogeneous media where $\mu$ and $\epsilon$ exhibit discontinuous behavior at media interfaces. Here, we present a solution of Maxwell’s equation in both a homogeneous and a heterogeneous medium via W-PINN and compare them with the traditional PINN.
Table 16: Parameters used for Example 3.8.
Parameters Value
Resolutions $(J_{x},J_{t})$ [-5,5], [-5,5]
Number of hidden layers 6
Neurons per layer 50
Number of collocation points $10^{4}$
Number of boundary points $500$
Number of initial points $500$
Figure 16: Top: $E_{y}$ predicted at the left and corresponding point-wise error at the right. Bottom: $H_{z}$ predicted and corresponding point-wise error for the Maxwell’s equation14 in homogeneous media via W-PINN.
Figure 17: Top: $E_{y}$ predicted at the left and corresponding point-wise error at the right. Bottom: $H_{z}$ predicted and corresponding point-wise error for the Maxwell’s equation 14 in heterogeneous media via W-PINN.
  1. A.
    Homogeneous media
We first investigate the performance of our proposed method by solving Maxwell’s equations in a one-dimensional cavity model with homogeneous media. The governing equations are:
$\frac{\partial E_{y}}{\partial t}=-\frac{1}{\epsilon}\frac{\partial H_{z}}{\partial x},~{}~{}~{}\frac{\partial H_{z}}{\partial t}=-\frac{1}{\mu}\frac{\partial E_{y}}{\partial x},\quad x\in[0,1],\quad t\in[0,1].$ (14)
The system is subject to perfectly electric conductor (PEC) boundary conditions:
$\begin{cases}E_{y}(0,t)=E_{y}(1,t)=0,\\ \dfrac{\partial H_{z}(x,t)}{\partial x}\bigg{|}_{x=0,1}=0.\end{cases}$ (15)
The ground truth for the example is given by:
$\begin{cases}E_{y}=\sin(n\pi x)\cos(\omega t),\\ H_{z}=-\cos(n\pi x)\sin(\omega t),\end{cases}$ (16)
where $n=4$ is the order of cavity mode, cavity frequency $\omega=n\pi/l$ and $l$ is the length of medium.
Over five random runs, W-PINN achieved average relative L2-errors of $5.74\pm 2.12\times 10^{-4}$ and $6.11\pm 3.01\times 10^{-4}$ for $E_{y}$ and $H_{z}$, respectively. In contrast, traditional PINNs exhibited notably higher errors of $3.08\pm 1.18\times 10^{-3}$ and $5.18\pm 1.93\times 10^{-3}$. The computational efficiency of W-PINN is evident in its training time, requiring an average of 4.2 minutes with 18 ms per iteration on Nvidia A6000 GPU. The network architecture parameters are tabulated in Table 16. Figure 16 presents the W-PINN predictions for both electric and magnetic fields, alongside their corresponding point-wise errors within the cavity. The remarkably low point-wise error distribution validates the high accuracy of our approach in capturing the electromagnetic field behavior.
  1. B.
    Heterogeneous media
For this case, we chose the Maxwell’s equation 14 in a heterogeneous medium with the following analytical solution:
$E_{y}=\begin{cases}\cos(2t-2x+1)\\ +0.5\cos(2t+2x-1),&\text{if }x\in\Omega_{1},\\[8.0pt] 1.5\cos(2t-3x+1.5),&\text{if }x\in\Omega_{2},\end{cases}$ (17)
$H_{z}=\begin{cases}\cos(2t-2x+1)\\ -0.5\cos(2t+2x-1),&\text{if }x\in\Omega_{1},\\[8.0pt] 0.5\cos(2t-3x+1.5),&\text{if }x\in\Omega_{2},\end{cases}$ (18)
where $\Omega_{1}=[0,0.5]$ and $\Omega_{2}=[0.5,1]$. Here $\mu=1,\epsilon=1$ in $\Omega_{1}$ and $\mu=4.5,\epsilon=0.5$ in $\Omega_{2}$. According to electromagnetic interface condition:
$\begin{cases}E_{y}(0.5,t)\big{|}_{x\in\Omega 1}=E_{y}(0.5,t)\big{|}_{x\in\Omega 2},\\[6.0pt] H_{z}(0.5,t)\big{|}_{x\in\Omega 1}=H_{z}(0.5,t)\big{|}_{x\in\Omega 2}.\end{cases}$ (19)
In this challenging context of heterogeneous media, W-PINN demonstrates remarkable accuracy, achieving average relative L2-errors of $2.41\pm 1.22\times 10^{-4}$ and $2.81\pm 1.51\times 10^{-4}$ for electric ($E_{y}$) and magnetic ($H_{z}$) fields, respectively. These results represent a two-order-of-magnitude improvement over traditional PINNs, which exhibit substantially higher errors of $1.37\pm 1.18\times 10^{-2}$ and $2.07\pm 0.51\times 10^{-2}$. W-PINN took an average of 6.1 minutes with around 30 ms/iteration on Nvidia A6000 GPU. The network parameters are the same as in Table 16. Figure 17 illustrates the W-PINN predictions for both electric and magnetic fields, alongside their corresponding point-wise errors. While the media interface presents the greatest computational challenge (point-wise error regions in red), W-PINN maintains impressive accuracy throughout the domain, successfully capturing the Electromagnetic field behavior across the media.
Table 17: Parameters used for Example 3.9.
Parameters Value
Resolutions $(J_{x},J_{t})$ [-10,5], [-10,5]
Number of hidden layers 10
Neurons per layer 80
Number of collocation points $2\times 10^{4}$
Number of boundary points $1000$
Figure 18: A lid-driven unit square cavity.
Figure 19: Predicted fluid speed for the steady-state lid-driven cavity flow 20 for Re=100 (left) and Re=400 (right) via W-PINN.
Figure 20: Comparison of W-PINN solution (dashed line) with Ghia’s benchmark (scatterd circles) for the steady-state lid-driven cavity flow 20 for Re=100 (left) and Re=400 (right).
Example 3.9.
Lid-Driven Cavity flow:
The steady two-dimensional incompressible lid-driven cavity flow is a classic benchmark problem in computational fluid dynamics. The flow is governed by the incompressible Navier-Stokes equations, which present significant challenges due to their nonlinear nature and the presence of pressure-velocity coupling. As illustrated in Figure 18, the problem consists of a unit square cavity where the top wall moves with a uniform velocity of 1 unit while maintaining no-slip conditions on all other walls. The governing Navier-Stokes equations for this system are:
$\left\{\begin{array}[]{ll}\dfrac{\partial u}{\partial x}+\dfrac{\partial v}{\partial y}=0,\\[10.0pt] u\dfrac{\partial u}{\partial x}+v\dfrac{\partial u}{\partial y}+\dfrac{\partial p}{\partial x}-\dfrac{1}{\text{Re}}\left(\dfrac{\partial^{2}u}{\partial x^{2}}+\dfrac{\partial^{2}u}{\partial y^{2}}\right)=0,\\[10.0pt] u\dfrac{\partial v}{\partial x}+v\dfrac{\partial v}{\partial y}+\dfrac{\partial p}{\partial y}-\dfrac{1}{\text{Re}}\left(\dfrac{\partial^{2}v}{\partial x^{2}}+\dfrac{\partial^{2}v}{\partial y^{2}}\right)=0,\end{array}\right.$ (20)
where $u$ and $v$ are the fluid velocities in $x$ and $y$ directions respectively, and $p$ is the fluid pressure. Re is Reynolds number which represents the ratio of inertial forces to viscous forces. We evaluate our W-PINN approach for two distinct flow regimes: Re = 100 and Re = 400, comparing our results against the widely-accepted benchmark solutions of Ghia et al.[50].
For this example, network’s parameters are tabulated in 18. Figure 19 presents the fluid flow speed distribution ($|V|=\sqrt{u^{2}+v^{2}}$) predicted by W-PINN, demonstrating successful capture of the characteristic vortex structure. Quantitative validation against Ghia’s benchmark data is illustrated in Figure 20. For Re = 100, the maximum absolute errors are $3.41\times 10^{-3}$ and $2.46\times 10^{-3}$ for $u$ and $v$ components respectively. At the higher Reynolds number of 400, where flow complexity increases, the errors remain within acceptable bounds at $4.79\times 10^{-2}$ and $6.69\times 10^{-2}$. These results demonstrate W-PINN’s robust capability in handling complex fluid dynamics problems to a good extent.
@@ 4 Conclusion
An efficient W-PINN for singularly perturbed problems is developed. This study successfully demonstrated the advantages of W-PINNs over other PINN-based methods in dealing with problems with steep gradients, rapid oscillations, and singular or multiscale behavior. The wavelet theory served as the foundation for the development of W-PINN, which combined the learning efficiency of PINN with the localization properties of wavelets (both in scale and space) to capture nonlinear information within localized regions. A limitation of the proposed method is that the number of required basis functions grows exponentially with dimension, leading to computational challenges for high-dimensional PDEs. Investigating dimensionality reduction or sparse representation techniques can address this limitation in future research. In addition to this, W-PINNs could prove effective for other problems that exhibit singular behavior, such as fractional differential models, which inherently possess such characteristics.
@@ Acknowledgments
Ratikanta Behera is supported by the Anusandhan National Research Foundation (ANRF), Government of India, under Grant No. EEQ/2022/001065. We would like to express our gratitude to the Editor for taking time to handle the manuscript and to anonymous referees whose constructive comments are very helpful for improving the quality of our paper.
Preprint
A preprint of this paper is available on arxiv with Ref. No. arXiv:2409.11847